OpenStudy (anonymous):

Let g(x) = x^2 + 3 and find a function f that produces the given composition.

OpenStudy (anonymous):

\[(f o g)(x)=x^{4}+6x^{2}+9\]

OpenStudy (anonymous):

I'm a bit confused at what the above compositon means is \[x^{4}+6x^{2}+9 \] f(x)?

OpenStudy (anonymous):

sorry I mean g(x)?

OpenStudy (debbieg):

You have \(g(x) = x^2 + 3 \) You want to "cook up" an f(x) such that you get: \((fog)(x)=x^4+6x^2+9\) Remember, \((fog)(x)\) is just \(f(g(x))\)

OpenStudy (debbieg):

So you need to exaluate \(f(x^2+3 )\), and get the result: \(x^4+6x^2+9\) Any ideas what choice for f(x) will accomplish that?

OpenStudy (debbieg):

In other words, you need to decide on a funcion f(x), such that: \(\Large f(x^2+3 )=x^4+6x^2+9\) Ask yourself, "what do I have to DO to \(\Large x^2+3 \) to get \(\Large x^4+6x^2+9\) ?

OpenStudy (anonymous):

You'd have to square x^2+3 or (x^2+3)^2

OpenStudy (debbieg):

Exactly... :) So THAT is what you want f(x) to do, to whatever is "inputted" into it. So, f(x) = ??

OpenStudy (anonymous):

x^2

OpenStudy (debbieg):

Tah-dah! :) Very good. Then: \(\Large f(x^2+3 )=(x^2+3 )^2=x^4+6x^2+9\)

OpenStudy (anonymous):

Wow, I didn't realize it was that easy. Thanks.